[Brahms-l] Shift report 20050610 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Fri Jun 10 2005 - 08:02:55 EDT
Shifter leader was Priyanka SAxena
Shifters were 


1)There was beam intially only for 2 hours.It was then dumped.They were trying to inject the beam but couldnt do due to power dip.At present they are trying to recover from power dip.<br>
2)The beam was only for 2 hours.Run:15354-15359<br>
3)Everything was fine except of the T3. T3_2 A_XU was tripping down frequently.I tried to turn it up but it remains up only for few minutes and again trippped.So I turn it off finally.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Jun 10 08:03:12 2005

This archive was generated by hypermail 2.1.8 : Fri Jun 10 2005 - 08:03:20 EDT